OpenStudy (anonymous):

two first steps in solving -3(8x - 12)= 36

OpenStudy (texaschic101):

-3(8x - 12) = 36 step 1 : distribute the -3 through the parenthesis -24x + 36 = 36 step 2 : subtract 36 from both sides -24x + 36 - 36 = 36 - 36 leaving you with : -24x = 0 x = 0